# Disentangling spin excitation continua in classical and quantum magnets using 2D nonlinear spectroscopy

### v1

@@ III.3 Classical Model: Finite Temperature
Next, we present the 2DCS for the pure Kitaev model at $T/|K|=0.001$, shown in Figs. 3F-H for $K=-1$ and Figs. 3L-N for $K=1$. Although the horizontal and vertical signals survive, the diagonal nonrephasing and antidiagonal rephasing signals are completely absent for both models, implying that these signals are extremely sensitive to thermal noise. The reasoning for the sensitivity to temperature in the classical model is as follows. At zero temperature, the spin configurations are in the lowest energy configuration (forming a degenerate classical ground state manifold), meaning that the spins are almost perfectly aligned towards the local effective field $\textbf{H}_{\text{eff}}=\frac{\partial H}{\partial\textbf{S}_{i}}$, which appears in Eq. 5. When performing the time evolution using the LLG equations, the right hand side of Eq. 5 is only finite when there are time-dependent pulse fields present, thus the only spin dynamics are due to the presence of the pulses. In contrast, as the temperature is increased, the spins become misaligned with the local field, meaning spin dynamics are already present due to thermal fluctuations even without the pump fields. While these thermally-induced fluctuations are essential for the emergent spin excitation continuum in the DSSF, they may wash out the coherent nonlinear dynamics induced by the pulsed fields in 2DCS. A more detailed discussion on the decoherence of the spectroscopy is found in Appendix B. On the other hand, the diagonal and antidiagonal signals may be more robust in the quantum model since the quantum coherent signals may be less sensitive to thermal fluctuations. For the simpler case of the 1D transverse-field Ising model, it was shown that such signals can survive at finite temperaturesWatanabe et al. (2024). Hence, the investigation of the effects of temperature on the 2DCS signals in the quantum Kitaev model is a fruitful topic for future studies.
@@ IV Spin excitation continua and 2D spectroscopy
Figure 5: Field dependence of the dynamical spin structure factor (A-D) and two-dimensional coherent spectroscopy (E-L) for $K=-1$, $\Gamma=0.25$, and $\Gamma^{\prime}=-0.02$. A-D are replotted from Zhang et al. (2023), and were performed at $T/|K|=0.001$. $|M_{NL}^{z}(\omega_{t},\omega_{\tau})|$ is shown for $T/|K|=0$ (E-H) and $T/|K|=0.001$ (I-L). Panel F includes black solid, dashed, and dotted lines at energies corresponding to the three high intensity peaks within the $\Gamma_{0}$ continuum in B. Panel G includes a solid black line at the energy corresponding to the high intensity peak in panel C at $\Gamma_{0}$. In panels H and L, TR=Terahertz Rectification, 2Q=2 Quantum, NR=Non-Rephasing, R=Rephasing, PP=Pump Probe. $\mathbf{M}$, The zero temperature classical phase diagram at each field. ZZ=Zig-Zag, 32=32-site, 50=50-site, 18=18-site, PM=Polarized paramagnet.
@@
Motivated by recent experiments using high magnetic fields to reveal a disordered quantum paramagnet at intermediate field strengthsZhou et al. (2023), we examine a minimal model of $\alpha$-RuCl3 at various magnetic fields. We choose the parameterization $(K,\Gamma,\Gamma^{\prime})=(-1,0.25,-0.02)$, which yields large unit cell magnetic orders at intermediate fields between the zig-zag (ZZ) order at zero field and polarized phase at high field strengths. From the DSSF computed in previous studies for this parameter set plotted in Fig. 5A-D, a continuum of excitations can be seen at intermediate fields in panels B (32-site unit cell) and C (50-site unit cell). At finite temperatures, these fields host competing large unit cell orders which form a thermal ensemble, giving rise to continuum-like behavior. Fig. 5C in particular bears resemblance to the continuum seen in the pure ferromangetic Kitaev model in Fig. 3A. From the DSSF alone, it is hard to distinguish between this scenario and the continuum arising from the fractionalized excitations in the Kitaev spin liquid.
@@
We present the 2DCS for the $K\Gamma\Gamma^{\prime}$ model at zero temperature in Figs. 5E-H and at finite temperature in Figs. 5I-L in the $z$ measurement channel. We observe sharp peaks at zero temperature appearing at energies corresponding to the coherent magnon branches at the $\Gamma_{0}$ point in the DSSF. Each peak corresponds to a distinct nonlinear optical process and will appear at frequencies with integer multiples of $E_{n}$, the energy of the $n^{\text{th}}$ magnon branch. These nonlinear processes are most clear in the polarized phase, where only the lower magnon branch contributes to the 2DCS response. For example, the peak at $(\omega_{t},\omega_{\tau})=(E_{\text{lower}},0)$ is the so-called “pump-probe” response (labelled in Figs. 5H and L by “PP”). In principle, each magnon band may exhibit such nonlinear signals, as described in detail in Lu et al. (2017), where they measured the 2DCS spectra for the isolated magnon modes of the canted antiferromagnet YFO. Thus, for larger unit cell orders, we expect the 2DCS to be a collection of nonlinear signals for each magnon mode. For the ZZ phase, we see a collection of peaks in panel E at energies corresponding to the lower magnon branch from panel A. The intensity of the upper magnon branch is much weaker than the lower one in the DSSF, thus the 2DCS is dominated by the latter. For the 32-site and 50-site orders, each excitation on the continuum could in principle contribute peaks corresponding to all of the nonlinear processes described above, along with sum and difference frequency generation between the modes. As a result, the 2DCS exhibits a dense collection of many discrete peaks. Although a continuum is present, we can make out peaks in the 2DCS at energy scales corresponding to the strongest intensity peaks in the DSSF. These energies are shown in F as the black solid, dashed, and dotted lines for the three intense branches from B, while the highest intensity peak from C is marked by black solid lines in G. This dense collection of peaks can be directly contrasted with the continuous linear signals seen in the pure Kitaev model in Fig. 3. Thus, although both scenarios yield similar excitation continua in the DSSF, 2DCS produces qualitatively different results, reflecting the fundamentally distinct physics underlying the broad spectra.
@@
At finite temperature, in general we observe far fewer features than at zero temperature, similar to the pure Kitaev model case. The disappearance of the features present at zero temperature is particularly dramatic in panels J through L. For example, only the PP peak remains for the polarized phase, and the 2Q, TR, NR, and R peaks all disappear. However, these peaks are still visible in experiments done at finite temperatureLu et al. (2017); Zhang et al. (2024a), whereas most of our peaks disappear at $T/|K|=0.001$. The fragility of the signals at finite temperature again demonstrates the sensitivity to thermal noise in our simulations, drawing attention to the role that quantum effects may play in stabilizing these signals in actual experiments.
@@ V Discussion
We also highlighted the usefulness of 2DCS in differentiating between distinct mechanisms that can give rise to spin excitation continua in INS. In the $K\Gamma\Gamma^{\prime}$ model, we previously reported a broad continuum at finite fields arising from the many excitations of an ensemble of large unit cell magnetic ordersZhang et al. (2023). Naively, this INS response looks similar to the one that arises from the classical Kitaev spin liquid state. However, using 2DCS, we observed a dense collection of discrete points in the nonlinear response of these large unit cell orders, in stark contrast with the sharp linear signals in the pure Kitaev model. Thus, while linear response fails to distinguish between excitation continua arising from spin liquid physics and large unit cell magnetic orders, 2DCS exhibits clear qualitative differences between the two cases.
Our work establishes the ability of 2DCS to probe properties of excitations in spin systems otherwise inaccessible by linear spectroscopies, not only for classical versus quantum phenomena, but also for revealing the intrinsic nature of the excitations hidden in the broad spin excitation continua seen in INS. However, a few open questions remain. For example, what happens to the spectra in the quantum model in the presence of an external magnetic field and finite temperature, especially in the crossover region between a KSL and the polarized paramagnet? This question is relevant for distinguishing between a putative intermediate-field spin liquid and the field-polarized state in candidate Kitaev materials. Away from the exactly solvable point, the breaking of flux conservation negates the strict flux selection rules and hard flux gap, posing a challenge in identifying distinguishing characteristics of quantum spin liquid physics. Additionally, in both the pure Kitaev and $K\Gamma\Gamma^{\prime}$ models, the 2DCS simulated using classical MD are very sensitive to thermal noise, as many of the characteristic excitations of the underlying ground state disappear. The robustness of these signals in experiments Lu et al. (2017); Zhang et al. (2024a) may imply the importance of quantum effects in stabilizing them in the presence of thermal fluctuations. More generally, clarifying the role of quantum effects in 2DCS is thus another important step in understanding the nonlinear dynamics of frustrated magnets, and for quantum materials in general.
@@ Appendix A Details of the numerical methods
Figure 6: Magnetization at zero versus finite temperature in the presence of pulse A at various delay times $\tau$ for $K=-1$. The shaded region depicts the response after the pulse arrived.
We used classical Monte Carlo (MC) techniques to obtain the spin configurations needed to compute the 2DCS. We implemented an adaptive Gaussian update methodAlzate-Cardona et al. (2019), where each random spin sampling process will attempt to update within a width $\sigma$ away from the current spin, where $\sigma$ gets updated every sweep to converge towards an “optimal” $\sigma$. This optimal $\sigma$ is such that the target acceptance rate is approximately 50%. We used this algorithm to study system sizes of up to $L=20\times 20\times 2$, and we thermalized 200 independent Markov chains to the desired temperature using $10^{6}$ MC sweeps via annealing. For faster convergence, we also used 100 overrelaxation sweeps for every MC sweep, where the spins are flipped along the local fieldsPixley and Young (2008). For the zero temperature ($T/|K|=10^{-7}$) configurations, we performed $10^{8}$ deterministic update sweeps, where we randomly aligned the spins towards the local fieldsJanssen et al. (2016).
The spin configurations are then used as initial configurations (IC) for molecular dynamics, where each measurement is time-evolved deterministically according to the Landau-Lifshitz-Gilbert equations of motion as described in the main textLakshmanan (2011). For the input pulses, we used Gaussian pulse shapes described by $A(t)=\exp{[(-t/2t_{0})^{2}]}\cos(2\pi f_{0}t)$ where $t_{0}=0.38|K|^{-1}$ and $f_{0}=0.33|K|$. The strengths of the A and B pulses were $0.1|K|S$. For the numerical integration, we used the SSPRK53 ODE solver from the DifferentialEquations.jl Julia packageRackauckas and Nie (2017); Zhang and Shu (2011). We used a time window of $t_{\max}=600|K|^{-1}$ in timesteps of $\delta t=0.25|K|^{-1}$ for the simulations. These results were then numerically Fourier transformed in the $t$ and $\tau$ dimensions to obtain the 2DCS spectra.

### v2

@@ III.3 Classical Model: Finite Temperature
Next, we present the 2DCS for the pure Kitaev model at $T/|K|=0.001$, shown in Figs. 3F-H for K=1K=-1 and Figs. 3L-N for $K=1$. Although the horizontal and vertical signals survive, the diagonal nonrephasing and antidiagonal rephasing signals are completely absent for both models, implying that these signals are extremely sensitive to thermal noise. The reasoning for the sensitivity to temperature in the classical model is as follows. At zero temperature, the spin configurations are in the lowest energy configuration (forming a degenerate classical ground state manifold), meaning that the spins are almost perfectly aligned towards the local effective field $\textbf{H}_{\text{eff}}=\frac{\partial H}{\partial\textbf{S}_{i}}$, which appears in Eq. 5. When performing the time evolution using the LLG equations, the right hand side of Eq. 5 is only finite when there are time-dependent pulse fields present, thus the only spin dynamics are due to the presence of the pulses. In contrast, as the temperature is increased, the spins become misaligned with the local field, meaning spin dynamics are already present due to thermal fluctuations even without the pump fields. While these thermally-induced fluctuations are essential for the emergent spin excitation continuum in the DSSF, they may wash out the coherent nonlinear dynamics induced by the pulsed fields in 2DCS. A more detailed discussion on the decoherence of the spectroscopy is found in Appendix B. On the other hand, the diagonal and antidiagonal signals may be more robust in the quantum model since the quantum coherent signals may be less sensitive to thermal fluctuations. For the simpler case of the 1D transverse-field Ising model, it was shown that such signals can survive at finite temperatureswatanabe_exploring_2024. Hence, the investigation of the effects of temperature on the 2DCS signals in the quantum Kitaev model is a fruitful topic for future studies.
@@ IV Spin excitation continua and 2D spectroscopy
Figure 5: Field dependence of the dynamical spin structure factor (A-D) and two-dimensional coherent spectroscopy (E-L) for K=1K=-1, $\Gamma=0.25$, and $\Gamma^{\prime}=-0.02$. A-D are replotted from zhang_spin_2023, and were performed at $T/|K|=0.001$. $|M_{NL}^{z}(\omega_{t},\omega_{\tau})|$ is shown for $T/|K|=0$ (E-H) and $T/|K|=0.001$ (I-L). Panel F includes black solid, dashed, and dotted lines at energies corresponding to the three high intensity peaks within the $\Gamma_{0}$ continuum in B. Panel G includes a solid black line at the energy corresponding to the high intensity peak in panel C at $\Gamma_{0}$. In panels H and L, TR=Terahertz Rectification, 2Q=2 Quantum, NR=Non-Rephasing, R=Rephasing, PP=Pump Probe. $\mathbf{M}$, The zero temperature classical phase diagram at each field. ZZ=Zig-Zag, 32=32-site, 50=50-site, 18=18-site, PM=Polarized paramagnet.
@@
Motivated by recent experiments using high magnetic fields to reveal a disordered quantum paramagnet at intermediate field strengthszhou_possible_2023, we examine a minimal model of $\alpha$-RuCl3 at various magnetic fields. We choose the parameterization $(K,\Gamma,\Gamma^{\prime})=(-1,0.25,-0.02)$, which yields large unit cell magnetic orders at intermediate fields between the zig-zag (ZZ) order at zero field and polarized phase at high field strengths. From the DSSF computed in previous studies for this parameter set plotted in Fig. 5A-D, a continuum of excitations can be seen at intermediate fields in panels B (32-site unit cell) and C (50-site unit cell). At finite temperatures, these fields host competing large unit cell orders which form a thermal ensemble, giving rise to continuum-like behaviorfranke_thermal_2022; zhang_spin_2023. Fig. 5C in particular bears resemblance to the continuum seen in the pure ferromangetic Kitaev model in Fig. 3A. From the DSSF alone, it is hard to distinguish between this scenario and the continuum arising from the fractionalized excitations in the Kitaev spin liquid.
@@
We present the 2DCS for the $K\Gamma\Gamma^{\prime}$ model at zero temperature in Figs. 5E-H and at finite temperature in Figs. 5I-L in the $z$ measurement channel. We observe sharp peaks at zero temperature appearing at energies corresponding to the coherent magnon branches at the $\Gamma_{0}$ point in the DSSF. Each peak corresponds to a distinct nonlinear optical process and will appear at frequencies with integer multiples of $E_{n}$, the energy of the $n^{\text{th}}$ magnon branch. These nonlinear processes are most clear in the polarized phase, where only the lower magnon branch contributes to the 2DCS response. For example, the peak at $(\omega_{t},\omega_{\tau})=(E_{\text{lower}},0)$ is the so-called “pump-probe” response (labelled in Figs. 5H and L by “PP”). In principle, each magnon band may exhibit such nonlinear signals, as described in detail in lu_coherent_2017, where they measured the 2DCS spectra for the isolated magnon modes of the canted antiferromagnet YFO. Thus, for larger unit cell orders, we expect the 2DCS to be a collection of nonlinear signals for each magnon mode. For the ZZ phase, we see a collection of peaks in panel E at energies corresponding to the lower magnon branch from panel A. The intensity of the upper magnon branch is much weaker than the lower one in the DSSF, thus the 2DCS is dominated by the latter. For the 32-site and 50-site orders, each excitation on the continuum could in principle contribute peaks corresponding to all of the nonlinear processes described above, along with sum and difference frequency generation between the modes. As a result, the 2DCS exhibits a dense collection of many discrete peaks. Although a continuum is present, we can make out peaks in the 2DCS at energy scales corresponding to the strongest intensity peaks in the DSSF. These energies are shown in F as the black solid, dashed, and dotted lines for the three intense branches from B, while the highest intensity peak from C is marked by black solid lines in G. This dense collection of peaks can be directly contrasted with the continuous linear signals seen in the pure Kitaev model in Fig. 3. Thus, although both scenarios yield similar excitation continua in the DSSF, 2DCS produces qualitatively different results, reflecting the fundamentally distinct physics underlying the broad spectra.
@@
At finite temperature, in general we observe far fewer features than at zero temperature, similar to the pure Kitaev model case. The disappearance of the features present at zero temperature is particularly dramatic in panels J through L. For example, only the PP peak remains for the polarized phase, and the 2Q, TR, NR, and R peaks all disappear. However, these peaks are still visible in experiments done at finite temperaturelu_coherent_2017; zhang_terahertz_2024, whereas most of our peaks disappear at $T/|K|=0.001$. The fragility of the signals at finite temperature again demonstrates the sensitivity to thermal noise in our simulations, drawing attention to the role that quantum effects may play in stabilizing these signals in actual experiments.
@@ V Discussion
We also highlighted the usefulness of 2DCS in differentiating between distinct mechanisms that can give rise to spin excitation continua in INS. In the $K\Gamma\Gamma^{\prime}$ model, we previously reported a broad continuum at finite fields arising from the many excitations of an ensemble of large unit cell magnetic orderszhang_spin_2023. Naively, this INS response looks similar to the one that arises from the classical Kitaev spin liquid state. However, using 2DCS, we observed a dense collection of discrete points in the nonlinear response of these large unit cell orders, in stark contrast with the sharp linear signals in the pure Kitaev model. Thus, while linear response fails to distinguish between excitation continua arising from spin liquid physics and large unit cell magnetic orders, 2DCS exhibits clear qualitative differences between the two cases.
Our work establishes the ability of 2DCS to probe properties of excitations in spin systems otherwise inaccessible by linear spectroscopies, not only for classical versus quantum phenomena, but also for revealing the intrinsic nature of the excitations hidden in the broad spin excitation continua seen in INS. However, a few open questions remain. For example, what happens to the spectra in the quantum model in the presence of an external magnetic field and finite temperature, especially in the crossover region between a KSL and the polarized paramagnet? This question is relevant for distinguishing between a putative intermediate-field spin liquid and the field-polarized state in candidate Kitaev materials. Away from the exactly solvable point, the breaking of flux conservation negates the strict flux selection rules and hard flux gap,meaning that the quantum response should more closely resemble the classical case. This poses a challenge in identifying distinguishing characteristics of quantum spin liquid physics. Additionally, in both the pure Kitaev and $K\Gamma\Gamma^{\prime}$ models, the 2DCS simulated using classical MD are very sensitive to thermal noise, as many of the characteristic excitations of the underlying ground state disappear. The robustness of these signals in experiments lu_coherent_2017; zhang_terahertz_2024 may imply the importance of quantum effects in stabilizing them in the presence of thermal fluctuations. More generally, clarifying the role of quantum effects in 2DCS is thus another important step in understanding the nonlinear dynamics of frustrated magnets, and for quantum materials in general.
@@ Appendix A Details of the numerical methods
Figure 6: Magnetization at zero versus finite temperature in the presence of pulse A at various delay times $\tau$ for K=1K=-1. The shaded region depicts the response after the pulse arrived.
We used classical Monte Carlo (MC) techniques to obtain the spin configurations needed to compute the 2DCS. We implemented an adaptive Gaussian update methodalzate-cardona_optimal_2019-1, where each random spin sampling process will attempt to update within a width $\sigma$ away from the current spin, where $\sigma$ gets updated every sweep to converge towards an “optimal” $\sigma$. This optimal $\sigma$ is such that the target acceptance rate is approximately 50%. We used this algorithm to study system sizes of up to $L=20\times 20\times 2$, and we thermalized 200 independent Markov chains to the desired temperature using $10^{6}$ MC sweeps via annealing. For faster convergence, we also used 100 overrelaxation sweeps for every MC sweep, where the spins are flipped along the local fieldspixley_large-scale_2008. For the zero temperature ($T/|K|=10^{-7}$) configurations, we performed $10^{8}$ deterministic update sweeps, where we randomly aligned the spins towards the local fieldsjanssen_honeycomb-lattice_2016.
The spin configurations are then used as initial configurations (IC) for molecular dynamics, where each measurement is time-evolved deterministically according to the Landau-Lifshitz-Gilbert equations of motion as described in the main textlakshmanan_fascinating_2011-1. For the input pulses, we used Gaussian pulse shapes described by $A(t)=\exp{[(-t/2t_{0})^{2}]}\cos(2\pi f_{0}t)$ where $t_{0}=0.38|K|^{-1}$ and $f_{0}=0.33|K|$. The strengths of the A and B pulses were $0.1|K|S$. For the numerical integration, we used the SSPRK53 ODE solver from the DifferentialEquations.jl Julia packagerackauckas_differentialequationsjl_2017-1; zhang_maximum-principle-satisfying_2011. We used a time window of $t_{\max}=600|K|^{-1}$ in timesteps of $\delta t=0.25|K|^{-1}$ for the simulations. These results were then numerically Fourier transformed in the $t$ and $\tau$ dimensions to obtain the 2DCS spectra.